# UUITP-12/25

### v2

@@ 1 Introduction
where $A$ is the area of the surface and $S_{\text{out}}$ is the entanglement entropy of the quantum fields in the exterior of the surface. The generalized second law reflects the thermal properties of the black hole horizon. Since no counterexamples to the generalized second law are known, Wall’s proof puts the singularity theorem on a firmer footing. Significant advancements in the last decade have further refined the singularity theorem, resulting in wider applicability (see [4] and references therein).
The study of quantum complexity has also led us to another version of the second law of black hole thermodynamics. The Second Law of Complexity (SLC) states that if the complexity of a black hole is below its saturation value—which is of the exponential order in the Bekenstein-Hawking entropy—it will always increase [5, 6]. The complexity of black holes can be holographically realized through the “Complexity=Volume” (CV) prescription, where complexity is defined as the volume of boundary-anchored extremal co-dimension one surface passing through the interior of the black hole. Therefore, it is natural to ask whether we can use the CV prescription in conjunction with the second law of complexity to arrive at a singularity theorem. In this paper, we show that the answer is affirmative.
The statement of SLC is expected to hold for any causal horizon (see Section 3 for explicit definitions). We do not prove it here; instead, we use it as input. The following heuristic picture explains the interplay between singularities and SLC. Consider a sheet of future-directed null rays in the interior of a black hole. Let us look at extremal surfaces confined to the region between these null rays. The volume of these surfaces decreases as we move it toward the singularity, as shown below:
@@ 2 Growth of Extremal Surfaces
We then specify the location of the null surfaces in Figure 2 by choosing their corresponding Eddington-Finkelstein coordinates to be a constant, which we will denote by $u_{0}$.
@@
We choose $\lambda$ to increase from left to right in the Penrose diagram. Evaluating the equations of motion at the turning point, we find that $p<0$.
@@
Here, $p^{2}$ plays the role of the total energy of the particle. The scattering picture is handy in characterizing the extremal surfaces since we do not have explicit closed-form expressions. The effective potential has a maximum at some $r=r_{a}$, which we refer to as the accumulation surface. The region around this surface plays a crucial role in the usual holographic volume complexity of the black hole [10, 11, 7, 9]. The effective potential goes to zero at $r=0$ and $r=r_{h}$.
@@
Now, let us read off the properties of the extremal curves using the effective potential. Consider the case in which the anchor point is near the bifurcation surface of the black hole, i.e., $r_{v}$ is close to $r_{h}$. The associated extremal surface is represented by curve $1$ in the left panel of Figure 3. In the scattering picture, the surface corresponds to a particle scattering off the right side of the potential (see curve 1 in the right panel of Figure 3). The particle starts at $r=r_{v}$, reaches a turning point at $r_{a}<r_{\min}<r_{v}$, and then returns to the anchor surface.
@@
We have used the symmetry about the $t=0$ surface to simplify the expression. Using the equations of motion, we get the expression
@@
The choice of sign is determined by the sign of $\dot{r}$. When $r_{\min}>r_{a}$, $\dot{r}$ is positive while it is negative for the extremal surfaces with $r_{\min}<r_{a}$. This reverses the limits of integration. In the scattering picture, this corresponds to choosing the opposite orientation of the particle. Using (2.5), we can express $v$ as another $r$ integral
@@
where we have used (2.7) to simplify the expressions. Using $r^{*}_{v}=(v-u_{0})/2$, we get
@@ 3 Geodesic Incompleteness from Second Law
Now let us define a notion of a causal horizon. Consider a future-infinite timelike worldline $W$. We define a future-causal horizon $H_{\text{fut}}$ as the boundary of the past of the worldline, that is, $H_{\text{fut}}=\partial I^{-}\left(W\right)$ [14]. We can verify that this definition includes black hole event horizons and observer-dependent horizons, such as Rindler and cosmological horizons.
Using the definition of a causal horizon, we state the second law as follows [5, 6]:
Second Law of Complexity (SLC): The volume of extremal surfaces anchored to a causal horizon $H_{\text{fut}}$ always increases as we move the anchor point along the surface.
It is important to note that we do not have a proof of SLC. Explicit computations of holographic complexity in various cases, for example, in static black holes [7], rotating black holes [15], even non-static black holes [16, 17] all point towards the validity of SLC. We will not attempt to prove SLC here. Instead, we assume its validity.
@@
In this section, we will work with globally hyperbolic spacetimes, where (a) there are no closed timelike curves, and (b) for any two points $p$ and $q$, $J^{+}(p)\cap J^{-}(q)$ is compact, where $J^{\pm}(p)$ is the causal future/past of $p$. Now, let us use these ingredients to arrive at a singularity theorem.
@@
Consider a globally hyperbolic spacetime $M$ with a trapped extremal surface. Then, the spacetime is null geodesically incomplete.
@@
Let us assume the trapped extremal surface is anchored to a null surface $N$. We will suppose for a contradiction that the affine parameter on $N$ can be extended infinitely into the future. This would then imply that $N$ is a causal horizon. From SLC, the volume of the trapped surface would always increase, contradicting the defining property of a trapped extremal surface. Therefore, the affine parameter on $N$ must terminate at a finite value.
Now, we can complete the proof using the original Penrose theorem [1] (see [3] for more details). The affine parameter on $N$ being finite makes it compact. Since $N\equiv\partial I^{+}(p)$ is achronal, the global hyperbolicity of the spacetime leads to null geodesic incompleteness. ∎
@@ 4.1 Reissner-Nordström Black Hole
where $r_{\pm}$ are the location of the horizons with $r_{+}>r_{-}$. The inner horizon is a Cauchy horizon and signals the breakdown of global hyperbolicity [12]. The original Penrose theorem and its entropic reformulations do not to apply to such black holes because of this breakdown of global hyperbolicity.
@@
In the scattering picture, these surfaces can be thought of as particles scattering off the effective potential $V_{\text{eff}}$ (see right panel of Figure 4). As we move the anchor point towards the singularity, the volume of extremal surfaces initially increases but begins to decrease once the anchor point crosses the accumulation surface. The volume growth rates are given by (2.13) and (2.14). The crucial difference from the uncharged case is that $p$ never goes to zero, even when the anchoring point has reached the singularity. From the equations of motion, we can see that $p$ vanishes only when the anchor point reaches the singularity at the $t=0$ slice (shown by the horizontal dotted line in the Penrose diagram in Figure 4). Therefore, the growth rate always remains positive. The fate of the trapped extremal surfaces is not as doomed as in the previous case, as their volumes never shrink to zero when the anchor point reaches the singularity. This is because the null anchoring surfaces go off to timelike $r=0$ slices in two different coordinate patches. This suggests that the vanishing of the volume of extremal surfaces ceases to be a suitable probe of the singularity.
@@
When the anchor point crosses the inner horizon, we find new behaviour. There are now two possible choices for the extremal surfaces. We have the usual spacelike surface (curve 2 in Figure 5). Additionally, we get a timelike surface that satisfies the same boundary conditions (curve 3 in Figure 5). The orientation of the scattering trajectory in Figure 5 is chosen to ensure that $\dot{r}$ remains real. The lack of a unique extremal surface is a consequence of the breakdown of global hyperbolicity of the spacetime. The growth rate of these timelike surfaces can be computed using the techniques in Section 2. The equations of motion of these surfaces can be obtained by performing the replacements $p\to ip$ and $\lambda\to i\lambda$ in (2.5). We can see that $p\geq 0$ from the new equations of motion. This corresponds to choosing $\lambda$ to increase in the downward direction in the Penrose diagram.
Figure 5: The teal curves correspond to spacelike extremal surfaces while the orange curve is timelike. The red dashed lines once again corresponds to the locations of the anchor surface.
As in the case of spacelike surfaces, there is an accumulation surface whose location is given by the minimum of the effective potential. We will denote the location of this surface by $r=\hat{r}_{a}$. The orientation of the particle in the scattering picture depends on its turning point. When $r_{\min}\geq\hat{r}_{a}$, the volume of these timelike extremal surfaces is given by
@@
More importantly, we observe that $p$ vanishes when $r_{v}=0$. In this case, the extremal surface simplifies and becomes the constant $r=0$ slice. Consequently, its growth rate also drops to zero. Since the singularity is the only region where the growth rate vanishes, this behavior can be used as a robust signature of the black hole singularity.
It is not entirely surprising that timelike extremal surfaces can sense timelike singularities. However, what timelike complexity corresponds to in a standard quantum mechanical system is not immediately apparent. Interestingly, timelike entanglement entropy has been proposed as a method for computing what is referred to as pseudo-entropy in the boundary theory [23, 24]. It is natural to expect a corresponding dictionary to exist in the case of complexity. We leave this for future work.
@@ 4.2 Kerr Black Hole
The black hole horizons are located at $r_{\pm}$. We can also define an Eddington-Finkelstein coordinate $v=t+r^{*}$, where
@@
As in the RN case, we study extremal surfaces by anchoring the ‘right’ endpoint of the extremal surfaces at a null surface in the interior of the black hole. The other endpoint is chosen to lie at its turning point. The spacetime is not globally hyperbolic, and as a result, when the anchor point crosses the inner horizon, there are new timelike extremal surfaces. As the anchor point approaches the singularity, the timelike extremal surfaces also approach the singularity. To see this, let us look at hypersurfaces of the form $r(\theta)$. If $r$ is a constant, the extremality of the corresponding volume element is given by the condition [22, 25]
@@
We can see that this is trivially satisfied by the singularity at $r=0$ and $\theta=\pi/2$. Therefore, the singularity is an extremal volume surface. Now, let us compute the growth rate of the timelike surfaces as they approach the singularity. The volume element of a constant $r$ hypersurface is given by [15]
$\mathrm{d}\mathcal{V}=\sqrt{\rho^{2}\Delta}\sin{\theta}\ \mathrm{d}t\wedge\mathrm{d}\theta\wedge\mathrm{d}\phi\,.$ (4.8)
Since $\mathrm{d}v=\mathrm{d}t$ on a constant $r$ slice, we have
@@
At the singularity, the integrand vanishes, and the growth of timelike extremal surfaces goes to zero. This suggests that the growth rate of timelike or spacelike extremal surfaces can serve as a useful probe of black hole singularities in a variety of cases.

### v3

@@ 1 Introduction
where $A$ is the area of the surface and $S_{\text{out}}$ is the entanglement entropy of the quantum fields in the exterior of the surface. The generalized second law reflects the thermal properties of the black hole. Since no counterexamples to the generalized second law are known, Wall’s proof puts the singularity theorem on a firmer footing. Significant advancements in the last decade have further refined the singularity theorem, resulting in wider applicability (see [4] and references therein).
The study of quantum complexity has also led us to another version of the second law of black hole thermodynamics. The Second Law of Complexity (SLC) states that if the complexity of a black hole is below its saturation value—which is of the exponential order in the Bekenstein-Hawking entropy—it will always increase [5, 6]. The complexity of black holes can be holographically realized through the “Complexity=Volume” (CV) prescription, where complexity is defined as the volume of boundary-anchored extremal co-dimension one surface passing through the interior of the black hole [7, 8]. Therefore, it is natural to ask whether we can use the CV prescription in conjunction with the second law of complexity to arrive at a singularity theorem. In this paper, we show that the answer is affirmative.
We begin by proposing a generalization of the second law of complexity: the volume of extremal surfaces anchored to a causal horizon increases as the anchor points are pushed to the future (see Section 3 for more details). While we do not provide a proof, we take this statement as an assumption, supported by various physically relevant examples. The second law of complexity provides a powerful bulk condition that can be used to establish a singularity theorem. The following heuristic picture explains the interplay between singularities and SLC. Consider a sheet of future-directed null rays in the interior of a black hole. Let us look at extremal volume surfaces confined to the region between these null rays. The volume of these surfaces decreases as we move it toward the singularity, as shown below:
@@ 2 Growth of Extremal Surfaces
We then specify the location of the null surfaces by choosing their corresponding Eddington-Finkelstein coordinates to be a constant, which we will denote by $u_{0}$.
@@
We choose $\lambda$ to increase from left to right in the Penrose diagram. Evaluating the equations of motion at the turning point, we find that $p$ is negative.
@@
Here, $p^{2}$ plays the role of the total energy of the particle. The scattering picture is handy in characterizing the extremal surfaces since we do not have explicit closed-form expressions. The effective potential has a maximum at some $r=r_{a}$, which we refer to as the accumulation surface. The region around this surface plays a crucial role in the usual holographic volume complexity computations [8, 7, 9, 11].
@@
Now, let us read off the properties of the extremal curves using the effective potential. Consider the case in which the anchor point is near the bifurcation surface of the black hole, i.e., $r_{v}$ is close to $r_{h}$. The associated extremal surface is represented by curve $1$ in the left panel of Figure 3. In the scattering picture, the surface corresponds to a particle scattering off the right side of the potential (see curve 1 in the right panel of Figure 3). The particle starts at $r=r_{v}$, reaches a turning point located at $r_{a}<r_{\min}<r_{v}$, and then returns to the anchor surface.
@@
We have used the symmetry about the $t=0$ surface to simplify the expression. Using the equations of motion, we get
@@
The choice of sign is determined by the sign of $\dot{r}$. When $r_{\min}\geq r_{a}$, $\dot{r}$ is positive while it is negative for the extremal surfaces with $r_{\min}<r_{a}$. This reverses the limits of integration. In the scattering picture, this corresponds to choosing the opposite orientation of the particle.
Using (2.5), we can express $v$ as another $r$ integral:
@@
We have used (2.10) to simplify the expression. Using $r^{*}_{v}=(v-u_{0})/2$, we get
@@ 3 Geodesic Incompleteness from Second Law
Now let us define a notion of a causal horizon. Consider a future-infinite timelike worldline $W$. We define a future-causal horizon $H_{\text{fut}}$ as the boundary of the past of the worldline, that is, $H_{\text{fut}}=\partial I^{-}\left(W\right)$ [14]. It is important to emphasize that $W$ is future-infinite; that is, the proper time along the worldline runs all the way to infinity. This distinguishes $H_{\text{fut}}$ from the future boundary of the domain of dependence of arbitrary subregions. We can then quickly verify that this definition of causal horizons includes black hole event horizons as well as observer-dependent horizons, such as Rindler and cosmological horizons.
Using the definition of a causal horizon, we rephrase the second law in [5, 6] as follows:
Second Law of Complexity (SLC): The volume of extremal surfaces anchored to a causal horizon $H_{\text{fut}}$ always increases as we move the anchor point along the surface into the future.
It is important to note that we do not have a proof of SLC. Explicit computations of holographic complexity in various cases, for example, in static black holes [9], Rindler wedges, rotating black holes [15], even non-static black holes [16, 17] all point towards the validity of SLC [^1] Readers familiar with holographic complexity computations might notice that the SLC used here differs from standard definition, as the extremal surface is anchored to the horizon rather than to a cutoff surface in the asymptotic region. However, we can quickly convince ourselves that the growth of these surfaces comes from the turning point, which lies well inside the horizon, and that both formulations are classically equivalent [9] (see also [8, 18, 11] for related discussions).. We will not attempt to prove SLC here. Instead, we assume its validity.
@@
In this section, we restrict ourselves to globally hyperbolic spacetimes, where (a) there are no closed timelike curves, and (b) for any two points $p$ and $q$, $J^{+}(p)\cap J^{-}(q)$ is compact, where $J^{\pm}(p)$ is the causal future/past of $p$.
The statement of the SLC turns out to be extremely powerful because it provides a bulk condition on the growth rate of volumes anchored to causal horizons. An immediate implication of the SLC is that if there are trapped extremal surfaces somewhere in the spacetime, then they cannot be anchored to a causal horizon. Now, let us use this fact to arrive at a singularity theorem.
@@
Consider a globally hyperbolic spacetime $M$ with a trapped extremal surface $T$. Then, the spacetime is null geodesically incomplete.
@@
Let us assume the trapped extremal surface $T$ is anchored to a null surface $N=\partial I^{+}(p)$, for some point $p\in M$. We will suppose for a contradiction that the affine parameter of every null geodesic generating $N$ can be extended infinitely into the future. This would then imply that $N$ is a causal horizon. From SLC, the volume of the trapped surface would always increase, contradicting the defining property of a trapped extremal surface. Therefore, the affine parameter of the null geodesics on $N$ must terminate at a finite value.
Now, we can complete the proof using the original Penrose theorem [1] (see [12, 3] for more details). Suppose $M$ were null geodesically complete. Then the null geodesic generators of $N$ could be extended to the future beyond $N$. An important point to note is that the affine parameter on $N=\partial I^{+}(p)$ is bounded, and any extension of the null generators to the future would lie inside $I^{+}(p)$, not on its boundary (see Proposition 4.5.14 and Section 8.2 of [12] for a related discussion). Null geodesic completeness of $M$ implies that the endpoints of the null geodesic generators of $N$ lie within the spacetime and therefore remain part of $N$, ensuring that $N$ is closed.
We now argue that $N$ is compact. Since the affine parameters are bounded, we can rescale them to lie within the range $[0,1]$. Then $N$ can be topologically written as $T\times[0,1]$, which is a product of a compact space and a closed interval, making $N$ compact.
This conclusion contradicts the assumption that the spacetime is globally hyperbolic. To see this, consider a connected, non-compact Cauchy slice $\Sigma$. We can always find a smooth timelike vector field $t^{\mu}$ whose integral curves intersect $\Sigma$ exactly once. Because $N\equiv\partial I^{+}(p)$ is achronal, these integral curves intersect $N$ at most once as well. This allows us to define a one-to-one map between $\Sigma$ and $N$ using the integral curves of $t^{\mu}$. However, this leads to a contradiction since the global hyperbolicity of the spacetime prohibits the evolution of a non-compact slice $\Sigma$ into a compact surface $N$. Therefore, the assumption that $M$ is null geodesically complete, which was used to argue that $N$ is compact, is incompatible with the other assumptions. ∎
Before we conclude the discussion, a few comments are in order:
@@
  • In the standard CV prescription, complexity is assigned to the volume of a surface that lies outside the accumulation surface. No such prescription exists for trapped extremal surfaces located inside the accumulation surface. While it may be possible to define a notion of complexity for these trapped surfaces, the proof does not rely on such a definition. The SLC is carefully phrased to depend only on the volume of extremal surfaces anchored to causal horizons. As a result, we avoid the need to interpret trapped extremal surfaces in terms of complexity and do not rely on any non-standard notions of holographic complexity.
@@ 4.1 Reissner-Nordström Black Hole
where $r_{\pm}$ are the location of the horizons with $r_{+}>r_{-}$. The inner horizon is a Cauchy horizon and signals the breakdown of global hyperbolicity [12]. The original Penrose theorem and its entropic reformulations do not apply to such black holes because of this breakdown of global hyperbolicity.
@@
In the scattering picture, these surfaces can be thought of as particles scattering off the effective potential $V_{\text{eff}}$ (see right panel of Figure 4). As we move the anchor point towards the singularity, the volume of extremal surfaces initially increases but begins to decrease once the anchor point crosses the accumulation surface. The volume growth rates are given by (2.13) and (2.14). The crucial difference from the uncharged case is that $p$ never goes to zero, even when the anchoring point has reached the singularity.
To see this, consider (2.13) where the growth rate can be seen to be proportional to $p$. From the first equation in (2.5), we observe that $p$ vanishes if and only if $f(r)=0$ or $t=$ constant along the extremal surface. Since the condition $f(r)=0$ is never satisfied, the only possibility is to consider null anchoring surfaces that reach the singularity at the $t=0$ slice (depicted by the horizontal dotted line in the Penrose diagram in Figure 4). For any other general anchoring surface, the growth rate remains nonzero. Therefore, the fate of the trapped extremal surfaces is not as doomed as in the previous case, as their volumes never shrink to zero when the anchor point reaches the singularity. This is because the null anchoring surfaces go off to timelike $r=0$ slices in two different coordinate patches. This suggests that the vanishing of the volume of extremal surfaces ceases to be a suitable probe of the singularity.
@@
When the anchor point crosses the inner horizon, we find new behaviour. There are now two possible choices for the extremal surfaces. We have the usual spacelike surface (curve 2 in Figure 5). Additionally, we get a timelike surface that satisfies the same boundary conditions (curve 3 in Figure 5). The orientation of the scattering trajectory in Figure 5 is chosen to ensure that $\dot{r}$ remains real. The lack of a unique extremal surface is a consequence of the breakdown of global hyperbolicity of the spacetime. The growth rate of these timelike surfaces can be computed using the techniques in Section 2. This time, we choose the parameter along the surface as follows:
$r^{d-1}\sqrt{-f\dot{v}^{2}+2\dot{v}\dot{r}}=-1\,,$ (4.2)
where the negative sign on the right-hand side comes from the fact that we are working with timelike surfaces. Denoting the conserved quantity along the surface by $p$, we arrive at the following equations of motion
[…]
Note that the equations of motion can also be obtained from (2.5) by the simple substitution $p\to ip$ and $\lambda\to i\lambda$. Choosing $\lambda$ to increase in the downward direction in the Penrose diagram, we find that $p\geq 0$.
Figure 5: The teal curves correspond to spacelike extremal surfaces, while the orange curve is timelike. The red dashed lines once again correspond to the locations of the anchor surface.
We can once again rewrite the equation involving the $r$ coordinate as a particle scattering of an effective potential $\tilde{V}_{\text{eff}}$. The new effective potential turns out to be the negative of the spacelike effective potential $V_{\text{eff}}$:
$\tilde{V}_{\text{eff}}(r)=-V_{\text{eff}}(r)=r^{2d-2}f(r)\,.$ (4.4)
As in the case of spacelike surfaces, there is an accumulation surface whose location corresponds to the maximum of $\tilde{V}_{\text{eff}}$, which is the same as the minimum of $V_{\text{eff}}$. We will denote the location of this surface by $r=\hat{r}_{a}$. The orientation of the particle in the scattering picture depends on its turning point w.r.t to the new accumulation surface. When $r_{\min}\geq\hat{r}_{a}$, the volume of these timelike extremal surfaces is given by
@@
More importantly, we observe that $p$ vanishes when $r_{v}=0$. In this case, the extremal surface simplifies and becomes the constant $r=0$ slice, and the right-hand side of the first equation of motion (4.3) vanishes identically. Consequently, the growth rate of the extremal surface drops to zero. Since the singularity is the only region where the growth rate vanishes, this behavior can be used as a robust signature of the black hole singularity.
While it is perhaps not surprising that timelike extremal surfaces are sensitive to timelike singularities, their dual interpretation remains unclear. In particular, the meaning of the volume of a timelike extremal surface in the boundary theory is not immediately evident. Nonetheless, it is tempting to speculate that this may hint at a notion of “timelike complexity” in the dual description. Interestingly, timelike entanglement entropy has been proposed as a method for computing what is referred to as pseudo-entropy in the boundary theory [24, 25, 26, 27]. It is natural to expect a corresponding dictionary to exist in the case of complexity. We leave this for future work.
@@ 4.2 Kerr Black Hole
The black hole horizons are located at $r_{\pm}$. The ring singularity sits at $r=0$ and $\theta=\pi/2$. We can also define an Eddington-Finkelstein coordinate $v=t+r^{*}$, where
@@
As in the RN case, we consider extremal surfaces anchored at a null surface inside the black hole. However, we now focus on axially symmetric configurations, as the black hole is rotating. We anchor the ‘right’ endpoint of the extremal surfaces at the null surface. The other endpoint is chosen to lie at its turning point. The spacetime is not globally hyperbolic, and as a result, when the anchor point crosses the inner horizon, there are new timelike extremal surfaces. In particular, as the anchor point approaches the singularity, the timelike extremal surfaces approach the singularity. Let us see how this works out.
Consider a general axisymmetric hypersurface in the interior of the black hole. The $r$ coordinate can be parametrized by $r(t,\theta)$ and the line element of this three-manifold can be computed by substituting $\mathrm{d}r=r_{,t}\mathrm{d}t+r_{,\theta}\mathrm{d}\theta$ into to the metric (4.7). Once we have the metric, we can compute the volume by integrating over the variables $t,\theta$, and $\phi$. The full expression is given in [23], although we will not need it explicitly here since we are only focusing on surfaces near the ring singularity.
We therefore restrict ourselves to the case where $r$ is just a function of $\theta$. We will also assume that the surface is $t$-independent. In this case, the three-metric computation simplifies significantly, and we obtain [15, 28]
$\mathrm{d}\mathcal{V}=\sqrt{\rho^{2}\left(|\Delta|-r_{,\theta}^{2}\right)}\sin\theta dt\wedge d\wedge d\theta\wedge d\phi\,.$ (4.10)
The extremal hypersurfaces are given by extremizing this integral over the volume element. If $r=$ constant, the Euler-Lagrange equation is given by the simple expression
@@
We can see that this is trivially satisfied by the ring singularity at $r=0$ and $\theta=\pi/2$. Therefore, the singularity is an extremal volume surface. Moreover, it is a timelike extremal surface.
Now, let us compute the growth rate of the timelike surfaces as they approach the singularity. Since $\mathrm{d}v=\mathrm{d}t$ on a constant $r$ slice, we can use the volume element (4.10) to obtain
@@
At the singularity, the integrand vanishes, and the growth of timelike extremal surfaces goes to zero. This suggests that the growth rate of timelike or spacelike extremal surfaces can serve as a useful probe of black hole singularities in various cases.